Please help~
If the equation kx^2+2x+28=k-x^2-6kx has equal roots, show that k=-9/10 or 3.
I tried substituting those numbers in but that didn't work out
Thank you.

Answer by ankor@dixie-net.com(22740)

the equation kx^2+2x+28=k-x^2-6kx has equal roots, show that k=-9/10 or 3.
kx^2 + 2x + 28 = k - x^2 - 6kx
replace k with 3
3x^2 + 2x + 28 = 3 - x^2 - 18x
Combine like terms on the left
3x^2 + x^2 + 2x + 18x + 28 - 3 = 0
4x^2 + 20x + 25 = 0
Factors to
(2x+5)(2x+5) = 0
2x = -5
x = -2.5, 2.5
:
I'll let you try k=-.9, it's past my bedtime